OpenStudy (anonymous):

father is five less than four times his son's age. eight years from now, he will be three more than twice his son's age. how old are they now?

OpenStudy (anonymous):

Father age = x Sons age = y 4y - 5 = x After 8 years x+8=2(y+8) + 3 4y - 5 +8 = 2y + 19 y = 8 x = 27